# Explainability of Text Processing and Retrieval Methods: A Survey

### v2

@@ 6.3. Global explanations: additive models as explainable surrogates or as inherently explainable models
Sen et al. (2020) studied simple, additive models as surrogates of both statistical and neural retrieval models, by projecting these models onto a 3-dimensional space, with well-understood dimensions, corresponding to term frequency (tf), document frequency (df) and document length (len). For each of the “complex” models, a simple, linear regression model was trained to predict the score of a term in a document using tf, idf and len. For neural models, semantic overlap was introduced as a fourth dimension to capture the contribution of soft (non-exact) term matches to the overall similarity between the query and a document. The regression models are intended to provide linear approximations (as shown in the example below) of the non-linear expressions used by ranking models:
@@
Generalised Additive Models (GAMs (Hastie and Tibshirani, 1986)) are, as the name suggests, a generalisation of linear regression models. Instead of learning just coefficients, a GAM learns a smooth function for each input feature. The outputs of these functions are simply added together to obtain the prediction $y$ for an input instance. GAMs are regarded as interpretable (Lou et al., 2012). Zhuang et al. (2021) applied GAMs to the Learning to Ranking (LETOR) problem. In their algorithm, NeuralRankGAM, each function constituting the GAM was represented by a simple feed forward neural network, leading to the claim that the resultant architecture is inherently interpretable. The authors adapted piece-wise regression (Muggeo, 2003) to construct simpler, distilled versions of these networks in order to make inferencing faster, as well as to make the models potentially more interpretable. Of course, it is debatable whether a “simple” model involving “only” a few thousand parameters is actually interpretable. On three LETOR datasets (Yahoo (Chapelle and Chang, 2011), Web30K (Qin and Liu, 2013), and a private dataset sub-sampled from the Chrome Web Store logs), NeuralRankGAM outperformed, in terms of NDCG, other GAMs (TreeGAM (Lou et al., 2012) and Tree RankGAM (Ke et al., 2017)) that were used as baselines. Interpretability was studied using visualisations of occlusion-based feature attribution scores (§ 4.3). The model appeared no more interpretable than its competitors.
Lucchese et al. (2022) proposed another “inherently interpretable” ranking model based on LambdaMART (Burges et al., 2011), that has a similar structure as GAMs, but uses ensembles of trees. Their approach uses a modified boosting LambdaMART algorithm in three stages. In the first stage, an ensemble of trees is used, with each tree working on a single feature. The interaction between distinct pairs of features (as identified in the first stage) is modeled in the second stage. In the last stage, the LambdaMART algorithm is restricted to use only the top $k$ feature pairs obtained from the previous stages. The model was compared against NeuralRankGAM (discussed above) and a pointwise ranking algorithm EBM (Lou et al., 2013), on three learning to rank datasets, Istella-S (Lucchese et al., 2016), Web30K, and Yahoo. The proposed model resulted in further improvements in NDCG, but its claim of intepretability is open to debate.
@@ 6.4. Global explanations based on probing
Recall that probing (Section 4.1) attempts to characterise linguistic features that are encoded within the representations generated at intermediate layers of a pre-trained network. In an IR context, early experiments with probing were reported by Palangi et al. (2016) and Cohen et al. (2018a). Cohen et al. trained a multilayered LSTM network for an IR-oriented task, viz., answer passage retrieval for the CQA (Surdeanu et al., 2008) and WikiQA (Yang et al., 2015) datasets. After the weights were frozen, the probing framework proposed in (Alain and Bengio, 2017) was applied to each layer of the network. The representations were used for NLP tasks such as POS tagging on the Penn TreeBank (Marcus et al., 1994), NER on CoNLL-2003 (Tjong Kim Sang and De Meulder, 2003), sentiment classification on the IMDB review collection (Maas et al., 2011), and textual entailment for Stanford’s Natural Language Inference (SNLI) corpus (Bowman et al., 2015). Cohen et al. found that initial layers appear to capture NER and POS features, while the later (second LSTM) layer is concerned with higher-level, abstract topical representations. Interestingly, the study by Palangi et al. indicates that such conclusions may be collection-dependent, i.e., the abstract representations appear to capture different features for different collections. For example, representations from the WikiQA model are more effective for POS tagging and NER than those from the CQA model, but they perform poorly on the sentiment classification and entailment tasks.
@@
Qiao et al. (2019) highlighted the interaction-based, sequence-to-sequence nature of BERT using a few different BERT-based rankers. A representation-based BERT model was designed by taking the individual [CLS] embeddings for both queries and documents, and computing the cosine similarity between them. Experiments on the MSMARCO passage reranking and TREC Web Track document ranking tasks showed that, on both datasets, the representation-based BERT model performed poorly. This suggests that BERT should not be regarded as a representation-based model. Interestingly, Qiao et al. observed that, while monoBERT (Nogueira and Cho, 2019) outperformed the then-state-of-the-art NRMs (K-NRM, Conv-KNRM) on MSMARCO, it failed to perform well on document ranking. They concluded that passage ranking for MSMARCO is more related to a question answering (QA) task, and therefore, better suited for a seq2seq model. Indeed, MSMARCO is basically a QA dataset, where precise matches are important. In contrast, traditional IR datasets, where recall is also important were used to train Conv-KNRM. It would be interesting to fine-tune BERT on such datasets, and then compare its performance against Conv-KNRM. On the other hand, their experiments showed that removing a single term from a document affects its score as calculated by the Conv-KNRM model, but monoBERT is more robust to such changes. Qiao et al. also contributed to the ongoing debate about attention (this is summarised in Section 9.1). They observed that the [CLS] and [SEP] tokens received the most attention. Further, even though removing stopwords does not affect the effectiveness of a ranker, stopwords and non-stop words received similar attention scores.
@@
A more recent series of experiments by MacAvaney et al. (2022) uses a different form of probing to analyse the sensitivity of NRMs to various changes in textual properties. In a sense, MacAvaney et al.’s approach is similar in spirit to feature-attribution studies: it looks at how the output changes when changes are made to input features. However, their objective is to obtain a global characterisation of NRMs, rather than determine which features of a particular input instance are important. A given NRM (trained on the MSMARCO passage ranking task) was first probed using a query-document pair $\left\langle Q,D\right\rangle$, and then using $\left\langle Q,D^{\prime}\right\rangle$, with $D^{\prime}$ being a perturbation of $D$. Four broad classes of perturbations were studied:  (a) “Measure and Match”, in which $D^{\prime}$ was generated by varying a specific attribute of $D$ (e.g. term frequency), while keeping other attributes (e.g., document length) fixed;  (b) simple textual manipulation (changing the order of words in a passage, introducing typographic errors, etc.);  (c) “Dataset Transfer Probes” (DTPs), which involve more complex, stylistic changes, e.g., changes in the level of fluency, formality, or the succinctness of a passage; and  (d) changes in factual content, which typically involve replacing one entity (e.g., New York) by another entity of the same type (e.g., Moscow). For the first two types of perturbations, synthetic documents were created, while DTPs were created from datasets specifically designed for the purpose (e.g., summarisation datasets). If a perturbation resulted in significantly different scores being computed for $D$ and $D^{\prime}$, the corresponding NRM was regarded as being sensitive to the perturbation. Experiments were conducted on the TREC-DL 19 passage ranking data and ANTIQUE (Hashemi et al., 2020). A wide range of models were studied: BM25, WMD (Kusner et al., 2015), SBERT (Reimers and Gurevych, 2019), LGBM (Ke et al., 2017), VBERT (Nogueira and Cho, 2019), T5 (Raffel et al., 2020), EPIC (MacAvaney et al., 2020), DT5Q (Nogueira and Lin, 2019), ANCE (Xiong et al., 2021b), and ColBERT (Khattab and Zaharia, 2020). Empirical observations include the following.
@@ 6.5. Global explanations for BERT-based NRMs: what knowledge do they possess?
Zhan et al. (2020) studied the monoBERT model focusing on both its attention mechanisms and query-document interactions. Once again, the MSMARCO passage ranking task was used for experiments. As in (Qiao et al., 2019) (discussed above) and similar studies conducted within the NLP community (Clark et al., 2019; Rogers et al., 2020; Kovaleva et al., 2019), empirical results show that BERT puts redundant attention weights on [CLS] and [SEP] tokens. Query-document interactions were found to play a major role in document ranking. Their observations regarding the behaviour of different layers of BERT echo earlier studies: interactions are not important in the initial layers of BERT, where it appears to learns more context specific information; in the later layers, it heavily uses interactions between a query and document to predict relevance. These findings were confirmed by experiments with probing (§ 4.1) and attribution (§ 4.3) techniques.
@@
$\mathit{Score}_{\mathrm{ColBERT}}(Q,D)=\sum_{q_{i}\in Q}\max_{d_{j}\in D}E_{q_{i}}E_{d_{j}}^{T}$
where $E_{q_{i}}$ and $E_{d_{j}}$ denote the embeddings of the $i$-th query term and $j$-th document token, respectively. This scoring function is superficially similar to early embedding-based similarity functions, except that $E_{q_{i}}$ and $E_{d_{j}}$ are heavily contextualised representations of query / document tokens. In earlier neural IR models, like DRMM (Guo et al., 2016), term heuristics (e.g., inverse document frequency) were explicitly involved in the architecture; ColBERT eliminates the need for such measures. Formal et al. (2021) present a “white-box analysis” of ColBERT on the TREC-DL 2019 and 2020 passage retrieval tasks. For each query term $q_{i}$, the authors analysed the distributions of exact matches and soft (non-exact) matches between $q_{i}$ and document tokens. They observed that exact matches are more important for high-IDF query terms. This is not unexpected, since such terms are likely to be specific, and therefore unlikely to have high semantic overlap with other terms in the document. The authors also estimated the impact of each $q_{i}$ on ColBERT’s ranking using a leave-one-out study: if the correlation ($\tau_{\mathit{AP}}$ (Yilmaz et al., 2008)) between the ranked lists for $Q$ and $Q\setminus\{q_{i}\}$ is high (close to 1, say), then $q_{i}$ is not important. Experiments show that ColBERT gives more importance to higher IDF terms, as expected.
@@
Fang et al. (2004) Proposed axiomatic framework TREC678 NA
Hagen et al. (2016) Relaxed axioms and used for re-ranking Clueweb09b github.com/webis-de/ir_axioms
Pang et al. (2017) Finding interaction vs representation focused NRMs Robust, LETOR 4.0 NA
Rennings et al. Diagnostic datasets with axioms Diagnostic dataset created from WikiPassageQA NA
Câmara and Hauff (2020) Diagnostic datasets with axioms Diagnostic dataset (proposed) github.com/ArthurCamara/bert-axioms
Sen et al. (2020) Additive models MSMARCO passage NA
Zhuang et al. (2021) GAMs EB30K, YAHOO, Chrome Web Store (private data) github.com/tensorflow/ranking/releases/tag/v0.3.1
Lucchese et al. (2022) Ensemble of trees WEB30K, YAHOO, Istella-S github.com/veneres/ilmart
Cohen et al. (2018a) Probing complex rankers CQA, WikiQA NA
Qiao et al. (2019) Understanding BERT for ranking MSMARCO passage, Clueweb09b NA
MacAvaney et al. (2022) Probing NRMs MSMARCO passage and created MMPs, TMPs from TREC-DL 2019 and MSMARCO passage github.com/allenai/abnirml
Formal et al. (2021) Understanding ColBERT TREC-DL 2019, 2020 passage NA

### v3

@@ 6.3. Global explanations: additive models as explainable surrogates or as inherently explainable models
10.1145/3397271.3401286 studied simple, additive models as surrogates of both statistical and neural retrieval models, by projecting these models onto a 3-dimensional space, with well-understood dimensions, corresponding to term frequency (tf), document frequency (df) and document length (len). For each of the “complex” models, a simple, linear regression model was trained to predict the score of a term in a document using tf, idf and len. For neural models, semantic overlap was introduced as a fourth dimension to capture the contribution of soft (non-exact) term matches to the overall similarity between the query and a document. The regression models are intended to provide linear approximations (as shown in the example below) of the non-linear expressions used by ranking models:
@@
Generalised Additive Models (GAMs (gam)) are, as the name suggests, a generalisation of linear regression models. Instead of learning just coefficients, a GAM learns a smooth function for each input feature. The outputs of these functions are simply added together to obtain the prediction $y$ for an input instance. GAMs are regarded as interpretable (10.1145/2339530.2339556). neural-rank-gam applied GAMs to the Learning to Ranking (LETOR) problem. In their algorithm, NeuralRankGAM, each function constituting the GAM was represented by a simple feed forward neural network, leading to the claim that the resultant architecture is inherently interpretable. The authors adapted piece-wise regression (est-regression-models-unk-break-points) to construct simpler, distilled versions of these networks in order to make inferencing faster, as well as to make the models potentially more interpretable. Of course, it is debatable whether a “simple” model involving “only” a few thousand parameters is actually interpretable. On three LETOR datasets (Yahoo (yahoo-ltr), Web30K (Qin2013IntroducingL4), and a private dataset sub-sampled from the Chrome Web Store logs), NeuralRankGAM outperformed, in terms of NDCG, other GAMs (TreeGAM (10.1145/2339530.2339556) and Tree RankGAM (lightgbm)) that were used as baselines. Interpretability was studied using visualisations of occlusion-based feature attribution scores (Section 4.2.2). The model appeared no more interpretable than its competitors.
ilmart proposed another “inherently interpretable” ranking model based on LambdaMART (pmlr-v14-burges11a), that has a similar structure as GAMs, but uses ensembles of trees. Their approach uses a modified boosting LambdaMART algorithm in three stages. In the first stage, an ensemble of trees is used, with each tree working on a single feature. The interaction between distinct pairs of features (as identified in the first stage) is modeled in the second stage. These features correspond to standard LETOR features, including BM25 score, document length, PageRank, etc. In the last stage, the LambdaMART algorithm is restricted to use only the top $k$ feature pairs obtained from the previous stages. The model was compared against NeuralRankGAM (discussed above) and a pointwise ranking algorithm EBM (ebm), on three learning to rank datasets, Istella-S (istella-s), Web30K, and Yahoo. The proposed model resulted in further improvements in NDCG, but its claim of intepretability is open to debate.
@@ 6.4. Global explanations based on probing
Probing (see also Section 4.1) attempts to characterise linguistic features that are encoded within the representations generated at intermediate layers of a pre-trained network. In an IR context, early experiments with probing were reported by deep-sentence-emb-for-lstm-to-ir and probing-bruce-croft-ictir-18. probing-bruce-croft-ictir-18 trained a multilayered LSTM network for an IR-oriented task, viz., answer passage retrieval for the CQA (surdeanu-etal-2008-learning) and WikiQA (yang-etal-2015-wikiqa) datasets. After training, the model weights were frozen, and the intermediate representations were subsequently used for probing—that is, to investigate what linguistic or semantic information was captured at different layers of the network, following a strategy similar to that proposed by alain2017understanding. At a high level, alain2017understanding employed linear classifiers to examine the extent to which internal representations within an image classification network encode specific forms of information. After training the network, they froze the weights and used the hidden-layer representations as features for simple linear classifiers to assess the presence of particular attributes. Building upon this idea, probing-bruce-croft-ictir-18 extended the probing methodology to several NLP tasks, including part-of-speech (POS) tagging on the Penn Treebank (penn-treebank), named entity recognition (NER) on CoNLL-2003 (conll), sentiment classification on the IMDB review dataset (maas-etal-2011-learning), and textual entailment using the Stanford Natural Language Inference (SNLI) corpus (snli-corpus). probing-bruce-croft-ictir-18 found that initial layers appear to capture NER and POS features, while the later (second LSTM) layer is concerned with higher-level, abstract topical representations. Interestingly, the study by deep-sentence-emb-for-lstm-to-ir indicates that such conclusions may be collection-dependent, i.e., the abstract representations appear to capture different features for different collections. For example, representations from the WikiQA model are more effective for POS tagging and NER than those from the CQA model, but they perform poorly on the sentiment classification and entailment tasks.
@@
Qiao2019UnderstandingTB highlighted the interaction-based, sequence-to-sequence nature of BERT using a few different BERT-based rankers. A representation-based BERT model was designed by taking the individual [CLS] embeddings for both queries and documents, and computing the cosine similarity between them. Experiments on the MSMARCO passage reranking and TREC Web Track document ranking tasks showed that, on both datasets, the representation-based BERT model performed poorly. This suggests that BERT should not be regarded as a representation-based model. Interestingly, Qiao2019UnderstandingTB observed that, while monoBERT (DBLP:journals/corr/abs-1901-04085) outperformed the then-state-of-the-art NRMs (K-NRM, Conv-KNRM) on MSMARCO, it failed to perform well on document ranking. They concluded that passage ranking for MSMARCO is more related to a question answering (QA) task, and therefore, better suited for a seq2seq model. Indeed, MSMARCO is basically a QA dataset, where precise matches are important. In contrast, traditional IR datasets, where recall is also important were used to train Conv-KNRM. It would be interesting to fine-tune BERT on such datasets, and then compare its performance against Conv-KNRM. On the other hand, their experiments showed that removing a single term from a document affects its score as calculated by the Conv-KNRM model, but monoBERT is more robust to such changes. Qiao2019UnderstandingTB also contributed to the ongoing debate about attention (this is summarised in Section D.1). They observed that the [CLS] and [SEP] tokens received the most attention. Further, even though removing stopwords does not affect the effectiveness of a ranker, stopwords and non-stop words received similar attention scores.
@@
A more recent series of experiments by 10.1162/tacl_a_00457 uses a different form of probing to analyse the sensitivity of NRMs to various changes in textual properties. In a sense, 10.1162/tacl_a_00457’s approach is similar in spirit to feature-attribution studies: it looks at how the output changes when changes are made to input features. However, their objective is to obtain a global characterisation of NRMs, rather than determine which features of a particular input instance are important. A given NRM (trained on the MSMARCO passage ranking task) was first probed using a query-document pair $\left\langle Q,D\right\rangle$, and then using $\left\langle Q,D^{\prime}\right\rangle$, with $D^{\prime}$ being a perturbation of $D$. Four broad classes of perturbations were studied:  (a) “Measure and Match”, in which $D^{\prime}$ was generated by varying a specific attribute of $D$ (e.g., term frequency), while keeping other attributes (e.g., document length) fixed;  (b) simple textual manipulation (changing the order of words in a passage, introducing typographic errors, etc.);  (c) “Dataset Transfer Probes” (DTPs), which involve more complex, stylistic changes, e.g., changes in the level of fluency, formality, or the succinctness of a passage; and  (d) changes in factual content, which typically involve replacing one entity (e.g., New York) by another entity of the same type (e.g., Moscow). For the first two types of perturbations, synthetic documents were created, while DTPs were created from datasets specifically designed for the purpose (e.g., summarisation datasets). If a perturbation resulted in significantly different scores for $D$ and $D^{\prime}$, the corresponding NRM was characterised as being sensitive to that perturbation. Experiments were conducted on the TREC-DL 19 passage ranking data and ANTIQUE (antique). A wide range of models were studied: BM25, WMD (word-mover-dist), SBERT (reimers-gurevych-2019-sentence), LGBM (lightgbm), VBERT (DBLP:journals/corr/abs-1901-04085), T5 (JMLR:v21:20-074), EPIC (epic), DT5Q (Cheriton2019FromDT), ANCE (ance), and ColBERT (colbert). Empirical observations include the following.
@@
We now discuss the probing framework used to analyse complex neural models, particularly large language models (LLMs). Low-Rank Adaptation (LoRA) (hu2022lora) has emerged as an efficient alternative to full fine-tuning of LLMs, offering comparable performance at a fraction of the computational cost. LoRA achieves this by freezing the pre-trained weights of an LLM and introducing trainable low-rank matrices into the Transformer architecture.
LoRA-based adaptation of LLMs for passage ranking on MS MARCO demonstrates promising retrieval performance. james-allan-short-probing examined how relevance (i.e., performance on retrieval tasks) is encoded across LLM layers, showing that LoRA updates to multi-head attention outperform those to multi-layer perceptron (MLP) layers, while jointly updating both further enhances ranking effectiveness across three LLMs: Llama 3.1–8B, Mistral–7B, and Pythia–6.9B. Notably, the study also found that mid-level MLP layers (layers 5–15 out of 32) contribute most to relevance prediction, emphasising the layer-specific nature of retrieval signals in LLMs. However, the reported ranking performance was evaluated only on the top-$10$ retrieved documents, whereas the average number of relevant documents per query in the TREC DL 2019 dataset is approximately $95$, suggesting a limited evaluation depth.
In a subsequent study, probing-llm-ictir-25 extended this line of work to analyse what types of features are encoded across different MLP layers. Specifically, they used four LLMs—Llama-2-7B, Llama-2-13B, Llama-3.1-8B, and Pythia-6.9B—each LoRA fine-tuned on the MS MARCO passage ranking dataset. Experimental results were reported on TREC DL 2019, TREC DL 2020, and the MS MARCO development query set. They examined a range of statistical features (e.g., average term frequency, maximum term frequency, BM25 score, and document length) alongside semantic features such as various distance-based query–document interaction measures. This analysis revealed that ranking-oriented LLMs also rely on term importance and exact matching signals, showing similar feature patterns across all four models. As in previous studies, while complex LLMs capture intricate relevance relationships, BERT and T5-based semantic features appear to be less prominently encoded. Interestingly, the analysis further showed that the final layer of the LLM encodes a linear combination of three simple features—query term frequency, variance of TF-IDF, and document length-normalised term frequency (tf)—indicating that a composite of these basic features plays a crucial role in modelling relevance.
@@ 6.5. Global explanations for BERT-based NRMs: what knowledge do they possess?
10.1145/3397271.3401325 studied the monoBERT model focusing on both its attention mechanisms and query-document interactions. Once again, the MSMARCO passage ranking task was used for experiments. As in (Qiao2019UnderstandingTB) (discussed above) and similar studies conducted within the NLP community (clark2019what; rogers-etal-2020-primer; kovaleva-etal-2019-revealing), empirical results show that BERT puts redundant attention weights on [CLS] and [SEP] tokens. Query-document interactions were found to play a major role in document ranking. Their observations regarding the behaviour of different layers of BERT echo earlier studies: interactions are not important in the initial layers of BERT, where it appears to learns more context specific information; in the later layers, it heavily uses interactions between a query and document to predict relevance. These findings were confirmed by experiments with probing (Section 4.1) and attribution (Section 4.2.2) techniques.
@@
(1) $\mathit{Score}_{\mathrm{ColBERT}}(Q,D)=\sum_{q_{i}\in Q}\max_{d_{j}\in D}E_{q_{i}}E_{d_{j}}^{T}$
where $E_{q_{i}}$ and $E_{d_{j}}$ denote the embeddings of the $i$-th query term and $j$-th document token, respectively. This scoring function is superficially similar to early embedding-based similarity functions, except that $E_{q_{i}}$ and $E_{d_{j}}$ are heavily contextualised representations of query / document tokens. In earlier neural IR models, like DRMM (drmm), term heuristics (e.g., inverse document frequency) were explicitly involved in the architecture; ColBERT eliminates the need for such measures. 10.1007/978-3-030-72240-1_23 present a “white-box analysis” of ColBERT on the TREC-DL 2019 and 2020 passage retrieval tasks. For each query term $q_{i}$, the authors analysed the distributions of exact matches and soft (non-exact) matches between $q_{i}$ and document tokens. They observed that exact matches are more important for high-IDF query terms. This is not unexpected, since such terms are likely to be specific, and therefore unlikely to have high semantic overlap with other terms in the document. The authors also estimated the impact of each $q_{i}$ on ColBERT’s ranking using a leave-one-out study: if the correlation ($\tau_{\mathit{AP}}$ (rank-correlation)) between the ranked lists for $Q$ and $Q\setminus\{q_{i}\}$ is high (close to 1, say), then $q_{i}$ is not important. Experiments show that ColBERT gives more importance to higher IDF terms, as expected.
@@
fang-sigir-04 Proposed axiomatic framework TREC678 NA
hagen-cikm-16 Relaxed axioms and used for re-ranking Clueweb09b github.com/webis-de/ir_axioms
Pang2017ADI Finding interaction vs representation focused NRMs Robust, LETOR 4.0 NA
rennings-diagnostic-dataset Diagnostic datasets with axioms Diagnostic dataset created from WikiPassageQA NA
camara-hauff-ecir Diagnostic datasets with axioms Diagnostic dataset (proposed) github.com/ArthurCamara/bert-axioms
10.1145/3397271.3401286 Additive models MSMARCO passage NA
neural-rank-gam GAMs EB30K, YAHOO, Chrome Web Store (private data) github.com/tensorflow/ranking/releases/tag/v0.3.1
ilmart Ensemble of trees WEB30K, YAHOO, Istella-S github.com/veneres/ilmart
probing-bruce-croft-ictir-18 Probing complex rankers CQA, WikiQA NA
Qiao2019UnderstandingTB Understanding BERT for ranking MSMARCO passage, Clueweb09b NA
10.1162/tacl_a_00457 Probing NRMs MSMARCO passage and created MMPs, TMPs from TREC-DL 2019 and MSMARCO passage github.com/allenai/abnirml
james-allan-short-probing Probing LLM-ranker MSMARCO passage NA
probing-llm-ictir-25 MSMARCO passage, TREC-DL 2019, 2020 github.com/TaKneeAa/ProbingRankLlama
10.1007/978-3-030-72240-1_23 Understanding ColBERT TREC-DL 2019, 2020 passage NA